OpenStudy (anonymous):

My math exam is tomorrow and its on Algebra I. Anyone wanna stick with me through some questions? (WILL FAN AND MEDAL ;)

OpenStudy (anonymous):

@alyssa_xo @dan815 @Legends @Michele_Laino @Nnesha @triciaal

OpenStudy (triciaal):

post the questions and someone will help

OpenStudy (anonymous):

how do i put these values: y = x - 1 and 2x + y = 11 in a table? (goes up by 1)

OpenStudy (anonymous):

Do you have anymore Q's?

OpenStudy (alyssa_xo):

solve for y, plot out your x's, plug in that's the standard procedure, I believe

OpenStudy (anonymous):

Yes, but the same type of questions.

OpenStudy (anonymous):

I would start off by putting both equations into slope-intercept form: y = mx + b ... which is practically solving for y.

OpenStudy (anonymous):

can u tell me if i get the right answer?

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

wait, so i just solve both the equations? using the substitution method?

OpenStudy (anonymous):

I'm not sure, and I don't want to give you an incorrect answer just in case. @alyssa_xo What do you think?

OpenStudy (dan815):

bibby lets talk on skype

OpenStudy (alyssa_xo):

bibowns, I'm in a call at the moment @Aureyliant , I'm not sure myself. is this a system of equations question or are you just plotting a table

OpenStudy (dan815):

sigh fine u want help on this stuff?? whats your question ask me ill teach you anything you want to learn

OpenStudy (anonymous):

^ My question exactly. Are you trying to combine both equations into just one table?

OpenStudy (dan815):

If you willing to learn I can make you ready, and get 100% on the text tomorrow

OpenStudy (dan815):

You have to do exactly as I say though

OpenStudy (dan815):

test*

OpenStudy (anonymous):

Wait dan are u talking to me lol? And yes @Aureyliant thats what i'm trying to do

OpenStudy (dan815):

yes I am

OpenStudy (anonymous):

Ok haha :) my question is at the top ^^^

OpenStudy (dan815):

im saying tell me what you want to learn

OpenStudy (dan815):

and ill teach u the stuff, not questions

OpenStudy (anonymous):

Okay, then yes, you'll need to solve for a system of equations using either graphically or algebraically.

OpenStudy (anonymous):

Oh how to put a system of equations in a table of values

OpenStudy (dan815):

huuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuh

OpenStudy (dan815):

to do what??

OpenStudy (dan815):

oh umm like this

OpenStudy (anonymous):

Once you have the combined system (or the result of the two equations) you can make a table of values based off of that

OpenStudy (anonymous):

so what would be the those two equations combined?

OpenStudy (anonymous):

and Dan do u know what ur talking bout

OpenStudy (anonymous):

Using the system of equations method.

OpenStudy (dan815):